Integration[-2xy/(x²+y²)²] with respect to y


1
Expert's answer
2021-01-29T12:56:10-0500

"\int \frac{-2xy}{(x^2+y^2)^2}dy=-x\int \frac{1}{(x^2+y^2)^2}d(y^2)=-x \int \frac{1}{(x^2+q)^2}dq" "=-x\int \frac{1}{w^2}dw=-x(- \frac{1}{w}+constant)" "=\frac{x}{w}+constant=\frac{x}{x^2+q}+constant" "=\frac{x}{x^2+y^2}+constant"